# Video Deepfake Abuse: How Company Choices Predictably Shape Misuse Patterns

### v1

@@ 1. Introduction
In 2022, AI image generation crossed a critical threshold, with models like DALL-E 2 (ramesh2021) and Stable Diffusion 1.x (Rombach2021) making the creation of photorealistic synthetic images much more efficient, dynamic, and accessible than ever before. In particular, Stable Diffusion’s 1.0’s open release in August 2022 made access to these capabilities widely available, allowing large numbers of users to create photorealistic images with minimal technical expertise, very little time, and no specialized data (petsiuk2022human). This enabled positive uses of image diffusion systems for recreation, graphics, and art. However, the accessibility of these new systems enabled unprecedented levels of misuse for AI-generated non-consensual intimate imagery (AIG-NCII), including AI-generated child sexual abuse material (AIG-CSAM). Since 2022, AIG-NCII and AIG-CSAM images have surged (securityhero2023; iwf2023ai-csam; iwf2024; NCMEC2024). For example, a report from ActiveFence estimated that the number of “threads related to the creation of [AIG-]NCII depicting private individuals rose…by 400%” between 2022 and 2023 (activefence2023). New image diffusion models (principally Stable Diffusion and Flux models, schneider2024image; hawkins2025deepfakes) and the infrastructure built around them (e.g., Chai2022) demonstrated how a small number of models capable of producing realistic NSFW deepfakes can drive misuse patterns (see Section 2).
@@
The image generation transformation of 2022 offers precedent for understanding current developments in video generation, where similar capability thresholds are now being crossed. Currently, in 2025, video generation models are undergoing an analogous revolution. Systems such as OpenAI’s Sora (openai2024), Google’s Veo (google2025), and Runway’s Gen-4 (runway2025) can be used to produce convincing videorealistic content. Meanwhile, as we will discuss in Section 3, a small number of open-weight models are emerging as the dominant tools for NSFW video generation, including Wan2.x. Variants of these models specialized for NSFW content are widely shared across several key online distribution platforms, including CivitAI, making these platforms critical gatekeepers for access to these capabilities.
@@ 2.1. Methods Prior to 2022
Prior to the rise of generative AI, convincing synthetic imagery depicting real humans was well-precedented. Early methods relied on manual image editing tools such as Photoshop, which could enable skilled users to stitch together images of different humans’ faces and bodies (eggestein2014fighting). In the 2010s, progress in research and development of generative AI added image and video processing models to the toolkit (monaghan2017impact; gieseke2020new; henry2020image). From 2017 onward, there was a notable rise in the availability of both “AI nudification” tools (for both images and video) and cultural awareness of them (kobriger2021out; wagner2019word). The gradual rise of tools that could enable AIG-NCII was driven, in part, by the development and release of open software tools from internet communities dedicated to synthetic NSFW content generation (monaghan2017impact; winter2020deepfakes).
Prior to 2022, despite it being possible to generate photorealistic NSFW images and videos of real individuals, tools were limited by skill requirements, a lack of dynamism, inefficiency, and/or inconsistent realism. For example, Photoshop required significant time and experience. Meanwhile, early generative adversarial networks (GANs) suffered from instability, mode collapse, and limited coherence, producing images that were often easily distinguishable from authentic photographs (saad2024). Prior to 2022, it was particularly challenging for GANs to consistently and dynamically generate realistic images. Models like AttnGAN and DM-GAN performed poorly on standard benchmarks and struggled with complex compositional understanding (xu2017; zhu2019). The most accessible approaches, such as VQGAN+CLIP combinations, required extensive technical expertise, complex prompt engineering, and produced imagery with characteristic visual artifacts that marked them as clearly synthetic (crowson2022; steinbruck2022). Even OpenAI’s DALL-E 1 diffusion model, while demonstrating impressive conceptual capabilities, operated at low resolution and was not publicly accessible (ramesh2021).
@@ 2.2. The DALL-E 2 Release
DALL-E 2 ushered in the modern age of dynamic photorealistic image generation. Released by OpenAI in April 2022, DALL-E 2 marked a leap forward in the ability of AI tools to consistently and dynamically produce approximately photorealistic results at high resolution (1024×1024) from simple text prompts (ramesh2022dalle2). OpenAI kept the model behind a closed API and reported implementing technical safeguards to prevent misuse, including filtering NSFW training data; internal testing; independent red team evaluations; prompt filtering to block explicit content and celebrity names; and filtering generated images for NSFW content. Furthermore, OpenAI required user registration with email verification and stated intentions; implemented user suspension systems for suspicious activity; and performed continuous backend updates to address emerging issues (dalle2modelcard). Subsequently, some researchers have found that these safeguards were imperfect, allowing for the production of some NSFW content (yang2024sneakyprompt). However, to the best of our knowledge, there is no evidence of successful, scalable uses of DALL-E 2 for the creation of individualized AIG-NCII.
@@ 2.3. Stable Diffusion’s Release and Downstream Misuse
Stability AI released Stable Diffusion 1.0 without effective mitigations against AIG-NCII misuse. Stable Diffusion 1.0 was released in August 2022 with open weights. It was released with a license, a user agreement, beta testing, and an NSFW classifier to block NSFW content. However, these measures proved to offer no substantial barrier to misuse. Usage terms are effectively unenforceable for an open-weight model; the beta testing did not reportedly involve red-teaming for AIG-NCII risks (stable_diffusion_launch_2022); and the NSFW content classifier could be trivially disabled by users who downloaded the system. [^2] The Stable Diffusion safety filter was also vulnerable to attacks (Rando2022). Meanwhile, Stable Diffusion 1.x models were trained on weakly-curated internet data (the LAION-5B dataset), which contained a substantial amount of NSFW content and CSAM, (birhane2311into; thiel2023). Meanwhile, Stability AI did not publicly report on any red teaming for NSFW capabilities or AIG-NCII risks. As a result of these ineffective safeguards, derivatives of Stable Diffusion 1.x models empirically became effective tools to generate AIG-NCII (schneider2024image; hawkins2025deepfakes).
@@
Community-developed tools greatly increased access to AIG-NCII creation in subsequent months. Despite the photorealistic image generation power of models like Stable Diffusion 1.x, they were not particularly dynamic or accessible tools by themselves. For example, initial misuse of them for AIG-NCII was largely focused on generating content of celebrities. The final key breakthrough toward widespread (mis)use was the introduction of software and applications that greatly increased dynamism and accessibility (wagner2025perpetuating). For example, DreamBooth, released in September 2022, enabled personalization of diffusion models using as few as three to five reference images (Chai2022). Using DreamBooth, users could “teach” models to associate unique identifiers with specific individuals, then generate novel synthetic content featuring those individuals in any described scenario. Meanwhile, other techniques, such as inpainting, outpainting, and stitching, [^3] Inpainting refers to having a diffusion model construct a customized synthetic version of an image region. Outpainting refers to having a diffusion model synthetically expand an image beyond its original border. Stitching refers to having a diffusion model realistically blend two separate images together (such as inserting one person’s head onto another person’s body). allowed users to produce NSFW content of specific subjects from SFW images (shani2021). These innovations meant that a person’s social media photos, school yearbook pictures, or family photos were sufficient for creating personalized photorealistic AIG-NCII images, expanding potential targeting from public figures to virtually anyone with minimal digital presence. Furthermore, community development quickly produced models further fine-tuned on NSFW data, user-friendly interfaces, pre-configured workflows, and optimized implementations accessible to non-technical users. This decentralized supply chain of these technologies creates a “malicious technical ecosystem” (ding2024malicious) that has enabled dozens of “AI nudifier” applications, which allow nontechnical users to create AIG-NCII within minutes (gibson2025analyzing; alexios_mantzarlis_ai_2025).
@@ 2.4. Misuse Patterns Following Technical Breakthroughs
A surge in personalized AIG-NCII followed the NSFW image generation breakthroughs of 2022. Within months of Stable Diffusion’s release, dedicated communities formed around generating non-consensual content, sharing techniques for circumventing safety measures, and distributing specialized model configurations optimized for explicit content generation. The vast majority of which depicts women and girls (securityhero2023; laffier2023deepfakes; my_image_my_choice_deepfake_2024; wagner2025perpetuating). Reports involving generative AI to the National Center for Missing and Exploited Children’s CyberTipline surged from 4,700 in 2023 to 67,000 in 2024, a 1,325% increase (NCMEC2024). [^4] These reports include AI-generated CSAM as well as other forms of child sexual exploitation involving generative AI, such as grooming guides and nudify applications. Analysis by Graphika found that 34 synthetic AIG-NCII providers received over 24 million unique visitors to their websites in September 2023, with referral link spam for these services increasing by more than 2,000% on platforms since the beginning of 2023. Graphika also found 52 Telegram groups used to access AIG-NCII services containing at least 1 million users as of September 2024 (graphika2023). Meanwhile, Activefence reported that “threads related to the creation of AIG-NCII depicting private individuals rose in the same period by 400%” (activefence2023). Ultimately, these developments have led to unprecedented patterns of AIG-NCII misuse. For example, in early 2025, law enforcement identified and made dozens of arrests connected to an international online community with hundreds of members dedicated to AIG-CSAM (europol2025operation).
@@ 3.1. Technical Capabilities and Accessibility
Video generation models in 2025 are crossing an analogous capability threshold to what enabled widespread AIG-NCII in 2022. Current systems, including Google’s Veo, OpenAI’s Sora, and Runway’s Gen-3, can often produce temporally coherent synthetic videos with quality approaching broadcast standards (openai2024; google2025; runway2025). Motion fidelity and facial consistency have reached levels where synthetic content sometimes becomes difficult to distinguish from authentic footage. [^5] Subjectively, we attest that a small fraction of the NSFW video deepfakes that we have seen online have appeared to us to be entirely indistinguishable from genuine video.
Model distribution platforms, principally CivitAI, facilitate the spread and usage of AI video models for NSFW content. Model distribution platforms provide infrastructure to help users download models and access detailed workflow documentation (gorwa2024moderating). By a large margin, the most prominent distribution platform for distributing models that specialize in NSFW content is CivitAI (wagner2025perpetuating; hawkins2025deepfakes; wei2024exploring; civitai2024transparency), who reported 446K LoRA adapters trained and 600M images/videos downloaded using their site in 2024. Community-developed workflows (e.g., from https://civitai.com/articles) further increase the accessibility of NSFW model generation capabilities by providing instructions, workflows, optimized settings, etc. that enable users to generate custom NSFW videos in minutes (fusionxcivitai2025).
@@ 3.2. Current Patterns and Methods for NSFW Video Generation
A small number of open-weight models account for the majority of model variants and videos across prominent online NSFW AI video communities To understand current usage patterns, we analyzed model search hits and tags across web communities. For this analysis, we assume that a model’s usage for generating NSFW content in general is an informative proxy for its usage in creating AIG-NCII (see Appendix A for further discussion). First, we identified 10 popular open-weight video generator model families online. [^6] We consider a model family to refer to a named and versioned set of models from a single developer, including all releases and derivatives. For example, we considered Wan-AI/Wan2.1-T2V-14B, Wan-AI/Wan2.2-TI2V-5B, and Comfy-Org/Wan_2.2_ComfyUI_Repackaged to all be from the same Wan2.x model family. We did this by exhaustively searching for all general-purpose text-to-video, image-to-video, and video-to-video models on Hugging Face from 2024 or later that had more than 10,000 downloads in the past month (as of October 8, 2025). Next, we identified a set of web platforms that allowed us to search and count AI video models specialized for SFW and NSFW content and AI videos containing SFW and NSFW content. Full details of our site selection process are in Appendix A. We ultimately selected six subreddits, CivitAI (the world’s most popular community platform for distributing AI models fine-tuned for generating NSFW content (civitai2024transparency)), and a website dedicated specifically to archiving models and content from CivitAI to evade moderation.
@@
As shown in Figure 2, different models’ SFW and NSFW market share differ across platforms. We order models based on their average NSFW/SFW market share ratio across all three analyses. Some models have disproportionately high market share for NSFW content. In particular, this includes Wan2.x, stable-video-diffusion, HunyuanVideo, and LTX-Video.  [^7] For comparison, we also analyzed xAI’s Grok Imagine model, which is fairly unique among frontier closed-weight AI video generators because its deployers have intentionally enabled its API to produce NSFW video content (maiberg2025grok; rainn2025grok). We analyzed the prevalence of videos from Grok Imagine on Reddit using the same methodology as with the open models. Despite accounting for 12.7% of NSFW discussions in our Reddit sample, Grok’s NSFW-to-SFW ratio was only 0.58, substantially lower than open-weight models like Wan2.x (1.39) and Stable Video Diffusion (1.04).

### v2

@@ 1. Introduction
In 2022, AI image generation crossed a critical threshold, with models like DALL-E 2 (Ramesh et al., 2021) and Stable Diffusion 1.x (Rombach and et al., 2021) making the creation of photorealistic synthetic images much more efficient, dynamic, and accessible than ever before. In particular, Stable Diffusion’s 1.0’s open release in August 2022 made access to these capabilities widely available, allowing large numbers of users to create photorealistic images with minimal technical expertise, very little time, and no specialized data (Petsiuk et al., 2022). This enabled positive uses of image diffusion systems for recreation, graphics, and art. However, the accessibility of these new systems enabled unprecedented levels of misuse for AI-generated non-consensual intimate imagery (AIG-NCII), including AI-generated child sexual abuse material (AIG-CSAM). Since 2022, AIG-NCII and AIG-CSAM images have surged (Hero, 2023; Internet Watch Foundation (IWF), 2023; Internet Watch Foundation, 2024; National Center for Missing & Exploited Children, 2024). For example, a report from ActiveFence estimated that the number of “threads related to the creation of [AIG-]NCII depicting private individuals rose…by 400%” between 2022 and 2023 (ActiveFence, 2023). New image diffusion models (principally Stable Diffusion and Flux models, Schneider and Hagendorff, 2024; Hawkins et al., 2025) and the infrastructure built around them (e.g., Chai and et al., 2022) demonstrated how a small number of models capable of producing realistic NSFW deepfakes can drive misuse patterns (see Section 2).
@@
The image generation transformation of 2022 offers precedent for understanding current developments in video generation, where similar capability thresholds are now being crossed. Currently, in 2025, video generation models are undergoing an analogous revolution (Internet Watch Foundation, 2026). Systems such as OpenAI’s Sora (OpenAI, 2024), Google’s Veo (Google, 2025), and Runway’s Gen-4 (Runway, 2025) can be used to produce convincing videorealistic content. Meanwhile, as we will discuss in Section 3, a small number of open-weight models are emerging as the dominant tools for NSFW video generation, including Wan2.x. Variants of these models specialized for NSFW content are widely shared across several key online distribution platforms, including CivitAI, making these platforms critical gatekeepers for access to these capabilities.
@@ 2.1. Methods Prior to 2022
Prior to the rise of generative AI, convincing synthetic imagery depicting real humans was well-precedented. Early methods relied on manual image editing tools such as Photoshop, which could enable skilled users to stitch together images of different humans’ faces and bodies (Eggestein and Knapp, 2014). In the 2010s, progress in research and development of generative AI added image and video processing models to the toolkit (Monaghan, 2017; Gieseke, 2020; Henry et al., 2020). From 2017 onward, there was a notable rise in the availability of both “AI nudification” tools (for both images and video) and cultural awareness of them (Kobriger et al., 2021; Wagner and Blewer, 2019). The gradual rise of tools that could enable AIG-NCII was driven, in part, by the development and release of open software tools from internet communities dedicated to synthetic NSFW content generation (Monaghan, 2017; Winter and Salter, 2020).
Prior to 2022, despite it being possible to generate photorealistic NSFW images and videos of real individuals, tools were limited by skill requirements, a lack of dynamism, inefficiency, and/or inconsistent realism. For example, Photoshop required significant time and experience. Meanwhile, early generative adversarial networks (GANs) suffered from instability, mode collapse, and limited coherence, producing images that were often easily distinguishable from authentic photographs (Saad and et. al., 2024). Prior to 2022, it was particularly challenging for GANs to consistently and dynamically generate realistic images. Models like AttnGAN and DM-GAN performed poorly on standard benchmarks and struggled with complex compositional understanding (Xu et al., 2017; Zhu et al., 2019). The most accessible approaches, such as VQGAN+CLIP combinations, required extensive technical expertise, complex prompt engineering, and produced imagery with characteristic visual artifacts that marked them as clearly synthetic (Crowson and Biderman, 2022; Steinbrück, 2022). Even OpenAI’s DALL-E 1 diffusion model, while demonstrating impressive conceptual capabilities, operated at low resolution and was not publicly accessible (Ramesh et al., 2021).
@@ 2.2. The DALL-E 2 Release
DALL-E 2 ushered in the modern age of dynamic photorealistic image generation. Released by OpenAI in April 2022, DALL-E 2 marked a leap forward in the ability of AI tools to consistently and dynamically produce approximately photorealistic results at high resolution (1024×1024) from simple text prompts (Ramesh et al., 2022). OpenAI kept the model behind a closed API and reported implementing technical safeguards to prevent misuse, including filtering NSFW training data; internal testing; independent red team evaluations; prompt filtering to block explicit content and celebrity names; and filtering generated images for NSFW content. Furthermore, OpenAI required user registration with email verification and stated intentions; implemented user suspension systems for suspicious activity; and performed continuous backend updates to address emerging issues (OpenAI, 2022). Subsequently, some researchers have found that these safeguards were imperfect, allowing for the production of some NSFW content (Yang et al., 2024a). However, to the best of our knowledge, there is no evidence of successful, scalable uses of DALL-E 2 for the creation of individualized AIG-NCII.
@@ 2.3. Stable Diffusion’s Release and Downstream Misuse
Stability AI released Stable Diffusion 1.0 without effective mitigations against AIG-NCII misuse. Stable Diffusion 1.0 was released in August 2022 with open weights. It was released with a license, a user agreement, beta testing, and an NSFW classifier to block NSFW content. However, these measures proved to offer no substantial barrier to misuse. Usage terms are effectively unenforceable for an open-weight model; the beta testing did not reportedly involve red-teaming for AIG-NCII risks (Lopez, 2022); and the NSFW content classifier could be trivially disabled by users who downloaded the system. [^2] The Stable Diffusion safety filter was also vulnerable to attacks (Rando and et al., 2022). Meanwhile, Stable Diffusion 1.x models were trained on weakly-curated internet data (the LAION-5B dataset), which contained a substantial amount of NSFW content and CSAM, (Birhane et al., 2023; Thiel, 2023). Meanwhile, Stability AI did not publicly report on any red teaming for NSFW capabilities or AIG-NCII risks. As a result of these ineffective safeguards, derivatives of Stable Diffusion 1.x models empirically became effective tools to generate AIG-NCII (Schneider and Hagendorff, 2024; Hawkins et al., 2025).
@@
Community-developed tools greatly increased access to AIG-NCII creation in subsequent months. Despite the photorealistic image generation power of models like Stable Diffusion 1.x, they were not particularly dynamic or accessible tools by themselves. For example, initial misuse of them for AIG-NCII was largely focused on generating content of celebrities. The final key breakthrough toward widespread (mis)use was the introduction of software and applications that greatly increased dynamism and accessibility (Wagner and Cetinic, 2025). For example, DreamBooth, released in September 2022, enabled personalization of diffusion models using as few as three to five reference images (Chai and et al., 2022). Using DreamBooth, users could “teach” models to associate unique identifiers with specific individuals, then generate novel synthetic content featuring those individuals in any described scenario. Meanwhile, other techniques, such as inpainting, outpainting, and stitching, [^3] Inpainting refers to having a diffusion model construct a customized synthetic version of an image region. Outpainting refers to having a diffusion model synthetically expand an image beyond its original border. Stitching refers to having a diffusion model realistically blend two separate images together (such as inserting one person’s head onto another person’s body). allowed users to produce NSFW content of specific subjects from SFW images (Shani and et al., 2021). These innovations meant that a person’s social media photos, school yearbook pictures, or family photos were sufficient for creating personalized photorealistic AIG-NCII images, expanding potential targeting from public figures to virtually anyone with minimal digital presence. Furthermore, community development quickly produced models further fine-tuned on NSFW data, user-friendly interfaces, pre-configured workflows, and optimized implementations accessible to non-technical users. This decentralized supply chain of these technologies creates a “malicious technical ecosystem” (Ding and Suresh, 2025) that has enabled dozens of “AI nudifier” applications, which allow nontechnical users to create AIG-NCII within minutes (Gibson et al., 2025; Alexios Mantzarlis and Santiago Lakatos, 2025).
@@ 2.4. Misuse Patterns Following Technical Breakthroughs
A surge in personalized AIG-NCII followed the NSFW image generation breakthroughs of 2022. Within months of Stable Diffusion’s release, dedicated communities formed around generating non-consensual content, sharing techniques for circumventing safety measures, and distributing specialized model configurations optimized for explicit content generation. The vast majority of which depicts women and girls (Hero, 2023; Laffier and Rehman, 2023; My Image My Choice, 2024; Wagner and Cetinic, 2025). Reports involving generative AI to the National Center for Missing and Exploited Children’s CyberTipline surged from 4,700 in 2023 to 67,000 in 2024, a 1,325% increase (National Center for Missing & Exploited Children, 2024). [^4] These reports include AI-generated CSAM as well as other forms of child sexual exploitation involving generative AI, such as grooming guides and nudify applications. Analysis by Graphika found that 34 synthetic AIG-NCII providers received over 24 million unique visitors to their websites in September 2023, with referral link spam for these services increasing by more than 2,000% on platforms since the beginning of 2023. Graphika also found 52 Telegram groups used to access AIG-NCII services containing at least 1 million users as of September 2024 (Lakatos, 2023). Meanwhile, Activefence reported that “threads related to the creation of AIG-NCII depicting private individuals rose in the same period by 400%” (ActiveFence, 2023). Ultimately, these developments have led to unprecedented patterns of AIG-NCII misuse. For example, in early 2025, law enforcement identified and made dozens of arrests connected to an international online community with hundreds of members dedicated to AIG-CSAM (Europol, 2025).
@@ 3.1. Technical Capabilities and Accessibility
Video generation models in 2025 are crossing an analogous capability threshold to what enabled widespread AIG-NCII in 2022. Current systems, including Google’s Veo, OpenAI’s Sora, and Runway’s Gen-3, can often produce temporally coherent synthetic videos with quality approaching broadcast standards (OpenAI, 2024; Google, 2025; Runway, 2025). Motion fidelity and facial consistency have reached levels where synthetic content sometimes becomes difficult to distinguish from authentic footage. [^5] Subjectively, we attest that a small fraction of the NSFW video deepfakes that we have seen online have appeared to us to be entirely indistinguishable from genuine video. These capabilities are now being used to create photorealistic AIG-CSAM at scale. In January 2026, the Internet Watch Foundation reported a “26,362% rise in [discovered] photo-realistic AI videos of child sexual abuse, often including real and recognisable child victims” between 2024 and 2025 (Internet Watch Foundation, 2026).
Model distribution platforms, principally CivitAI, facilitate the spread and usage of AI video models for NSFW content. Model distribution platforms provide infrastructure to help users download models and access detailed workflow documentation (Gorwa and Veale, 2024). By a large margin, the most prominent distribution platform for distributing models that specialize in NSFW content is CivitAI (Wagner and Cetinic, 2025; Hawkins et al., 2025; Wei et al., 2024; Civitai, 2024), who reported 446 thousand LoRA adapters trained and 600M images/videos downloaded using their site in 2024. Community-developed workflows (e.g., from https://civitai.com/articles) further increase the accessibility of NSFW model generation capabilities by providing instructions, workflows, optimized settings, etc. that enable users to generate custom NSFW videos in minutes (User, 2025).
@@ 3.2. Current Patterns and Methods for NSFW Video Generation
A small number of open-weight models account for the majority of model variants and videos across prominent online NSFW AI video communities To understand current usage patterns, we analyzed model search hits and tags across web communities. For this analysis, we assume that a model’s usage for generating NSFW content in general is an informative proxy for its usage in creating AIG-NCII (see Appendix A for further discussion). First, we identified 10 popular open-weight video generator model families online. [^6] We consider a model family to refer to a named and versioned set of models from a single developer, including all releases and derivatives. For example, we considered Wan-AI/Wan2.1-T2V-14B, Wan-AI/Wan2.2-TI2V-5B, and Comfy-Org/Wan_2.2_ComfyUI_Repackaged to all be from the same Wan2.x model family. We did this by exhaustively searching for all general-purpose text-to-video, image-to-video, and video-to-video models on Hugging Face from 2024 or later that had more than 10,000 downloads in the past month (as of October 8, 2025). Next, we identified a set of web platforms that allowed us to search and count AI video models specialized for SFW and NSFW content and AI videos containing SFW and NSFW content. Full details of our site selection process are in Appendix A. We ultimately selected six subreddits, CivitAI (the world’s most popular community platform for distributing AI models fine-tuned for generating NSFW content (Civitai, 2024)), and a website dedicated specifically to archiving models and content from CivitAI to evade moderation.
@@
As shown in Figure 2, different models’ SFW and NSFW market share differ across platforms. We order models based on their average NSFW/SFW market share ratio across all three analyses. Some models have disproportionately high market share for NSFW content. In particular, this includes Wan2.x, stable-video-diffusion, HunyuanVideo, and LTX-Video.  [^7] For comparison, we also analyzed xAI’s Grok Imagine model, which is fairly unique among frontier closed-weight AI video generators because its deployers have intentionally enabled its API to produce NSFW video content (Maiberg, 2025c; RAINN, 2025). We analyzed the prevalence of videos from Grok Imagine on Reddit using the same methodology as with the open models. Despite accounting for 12.7% of NSFW discussions in our Reddit sample, Grok’s NSFW-to-SFW ratio was only 0.58, substantially lower than open-weight models like Wan2.x (1.39) and Stable Video Diffusion (1.04). Note: Our data collection occurred in November 2025. In late December 2025, after our analysis period, Grok’s image editing feature was widely misused to create non-consensual intimate imagery, generating up to 7,751 sexualized images per hour at its peak (Collier et al., 2026). This demonstrates that even closed-weight models can face severe misuse when safety controls are insufficient.